# Algebraic geometry of bubbling Kähler metrics

### v1

@@ 2.2.1. Setup and algebraic construction
Hence, we can apply semistable reduction type theorem [Od24b, Theorem 3.8 (also cf., 1.1)] (which generalizes [Lan75, AHLH23]) and we obtain an Langton or [AHLH23, 3.35] type modification of the original morphism $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to[M/T]$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$) which corresponds to $\mathcal{X}$. We denote the obtained family as $\mathcal{X}^{\prime}_{\rm min,0}$ and going to confirm the required conditions in our Theorem 2.4. As noted in [Od24b], note that the family is non-canonical which corresponds to a choice of $T^{\prime}$ while the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ does not depend on it, from our construction (and [Od24b, §3]). Here, $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ denotes the discrete valuation ring which consists of the convergent power series of $t$ (for some positive unspecified radius). The obtained map from $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), or its finite extension $\mathbb{C}[[t^{\frac{1}{d}}]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{\frac{1}{d}}]]$) for some $d\in\mathbb{Z}_{>0}$, gives the desired family $\mathcal{X}^{\prime}_{\rm min}\to\Delta$ from the construction. The K-semistability of $\mathcal{X}^{\prime}_{\rm min,0}$ in the sense of [Sun23, §5] follows from the construction as $W$ is a K-semistable log terminal cone ($\mathbb{Q}$-Fano cone) in the sense of [CS18, CS19]. Furthermore, from the positivity of the weights on the coordinates of $T^{\prime}$-action and the definition of elementary modification [AHLH23, §3], we conclude that the obtained test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ is a negative weight deformation in the sense of [CH22]111111just of rank $1$ i.e., with constant $\xi_{i}$ in their definition, [Od24b, §2]. Note that we do not know yet if any good choice of $T^{\prime}$ gives rise to a $\Theta$-strata in the original sense, and independence of $T^{\prime}$ is a priori unclear (without using [Od24b]) even if that is the case. These points are why we need the higher rank setup and arguments in the generalized form [Od24b, Theorem 3.8] rather than the original in [AHLH23, §6].
@@
Now we are ready to discuss algebraic construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$, which follows a similar line of ideas as those for $\mathcal{X}^{\prime}_{\rm min}$, but also with some differences. We take a (non-canonical) algebraic subtorus $T^{\prime\prime}$ of $T$, with the rank $1$, which degenerates $W$ to the K-polystable degeneration as in [LWX21, Theorem 1.2]. Denote the corresponding base i.e., the partial compactification of $T^{\prime\prime}$ as $\overline{T}^{\prime\prime}$ which is isomorphic to $\mathbb{A}^{1}$. Recall that the construction in loc.cit is far from canonical. We still take one of them and denote the corresponding test configuration as $\mathcal{T}^{\prime\prime}\to\overline{T}^{\prime\prime}$. Henceforth we can and do assume $\mathcal{T}^{\prime\prime}$ is also achieved inside $M$, by enlarging $M$ inside the original base $\mathbb{A}_{k}^{m}$ if necessary, and $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) by retaking larger $l$, $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) and the embedding $\mathcal{X}\hookrightarrow\mathbb{C}^{l}\times\Delta$ (resp., $\mathcal{X}\hookrightarrow\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) is also realized in $M$ if necessary. Now we consider a couple of morphisms $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to M$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to M$) corresponding to $\mathcal{T}^{\prime}$ of degeneration of $X^{\prime}\rightsquigarrow W$, and corresponding to $\mathcal{T}^{\prime\prime}$ whose degeneration is of the form $W\rightsquigarrow C$.
@@
The above replacement $\mathcal{X}$ by $\mathcal{X}^{\prime}$ or $\mathcal{X}^{\prime\prime}$ or the existence of bubbles can first look in a sharp contrast to what happens in analogous global setup i.e., the separatedness of the moduli of log terminal projective Calabi-Yau varieties ([Od12, 4.2 (i), 4.3],  [^12] However, although 4.3 is fine, there is an obvious typo in its 4.2(i), mixing up notation with canonically polarized case: in the left hand side $-K_{\mathcal{X}/\mathcal{C}}$ should be read $\mathcal{L}$ and we assume K-stability (log terminality) of the fibers of $\pi$. [Bou14, 2.1]).
@@
Our proof follows that of Theorem 2.4 closely. For $\pi^{-1}(0)=\tilde{\mathcal{X}}_{0}$, we apply the same arguments as the beginning of the proof of Theorem 2.12. In particular, we take the same log terminal family $\mathcal{U}\subset\mathbb{A}_{\mathbbm{k}}^{l}\times\mathbb{A}_{\mathbbm{k}}^{m}$, which restricts to $\mathcal{V}\to M$.
We set $\mathcal{M}:=[M/T]$ again and consider the map $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to M\to\mathcal{M}$ which corresponds to $\tilde{\pi}$. As in the previous proof of Theorem 2.4, take the algebraic torus $T$ for $W$, the semistable tangent cone of $X$ and a rational polyhedral cone $\tau$ in the cocharacter lattice $N:={\rm Hom}(\mathbb{G}_{m},T)$ with $\xi\in\tau$ such that the base change of $\mathcal{V}$ over $U_{\tau}$ gives a flat isotrivial degeneration (over possibly higher dimensional base) with the central fiber $W$. If there is a uniform family of such isotrivial degenerations $X\rightsquigarrow W$ over a connected base, then from the same reason as in the proof of Theorem 2.4, i.e., by the use of index character ([MSY08, CS18, CS19]), the normalized volume are constant. Similarly as before, combined with the uniqueness of the $\widehat{\rm vol}$-minimizing valuation, this gives the canonical structure of higher $\Theta_{\tau}$-strata on $M_{c}$ for $c:=\widehat{\rm vol}(X,0)$. Then, we again apply the higher $\Theta$-stable reduction theorem [Od24b, §3.2, Theorem 3.8] to construct $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$. From the above discussions, this construction is completely canonical. The rest of the constructions are almost the same, so we omit the essential repetition. ∎
@@ 2.2.2. Comparison with the differential geometric construction
Now, we discuss comparison with the differential geometric construction in [DS14, Sun23] and prove Theorem 2.8.
@@
We prove the item ($\mathrm{i}$) of Theorem 2.8. In this proof, we closely follow the arguments and notations of [Sun23] (cf., also [DS17]) by fitting it to our notation in Theorem 2.8. In the notation of loc.cit, suppose that his $X_{j}(j=1,2,\cdots)$ are taken as fibers of our $\pi$ i.e., $X_{j}=X_{s_{j}}$ for some sequence $s_{j}\in\Delta$ with $s_{j}\to 0$ for $j=1,2,\cdots$ and $X_{\infty}=\pi^{-1}(0)=X$. Our $\xi$ corresponds to the real $N\times N$ diagonal matrix $\Lambda={\rm diag}(2^{\zeta_{1}},\cdots,2^{\zeta_{N}})$ with $\zeta_{i}\in\mathbb{R}_{>0}$s acting on the ambient space $\mathbb{C}^{N}$ in his notation. We crucially use the local holomorphic functions $G^{i},S^{i}_{j},L^{i}_{j}$ (and $H^{i}$, $H^{i}_{j}$ with higher degrees with respect to $\xi$) on the ambient spaces $\mathbb{C}^{N}$ of $X_{j}$s ($i$ was denoted as $b$ in loc.cit), $X_{\infty}$ and the possible bubbles $Z$, the bounds of rescaling orders $\underline{k}$ and $\overline{k_{j}}$ as prepared in loc.cit. Then, we replace the sequence $\{j\}$ by its subsequence so that there is a sequence of real numbers $b_{j}(j=1,2,\cdots)$ with $b_{j}\to\infty$ obtain that
$W_{1}:$ $=\lim_{j\to\infty}\Lambda^{b_{j}}\cdot X_{j}$
Let us write $\operatorname{deg}_{\xi}$ for the degree with respect to $\xi$ as in [Sun23, §4]. Here, each $L^{i}_{\infty}$ is obtained as a rescaled limit of $L^{i}_{j}$ for $j\to\infty$. Indeed, it can be achieved as follows. For each $i$, consider the set $L(i)$ of monomials $\chi^{\vec{m}}$ on the ambient space $\mathbb{C}^{N}$ whose $\xi$-degree is less than that of $G^{i}$. Its order is finite as $\zeta_{1},\cdots,\zeta_{N}$ are all positive. We denote the coefficient of $\chi^{\vec{m}}$ in $L^{i}_{j}$ as $c^{i}_{j}(\vec{m})$, then of course we have $c^{i}_{j}(\vec{m})\to 0\hskip 11.38109pt(j\to\infty)$ as $L^{i}_{j}\to 0\hskip 11.38109pt(j\to\infty)$. Now, for each fixed $j$, we consider
$b_{j}:=\max_{i,\vec{m}\in L(i)}\dfrac{\log(|c^{i}_{j}(\vec{m})|)}{\log(2)\cdot(\operatorname{deg}_{\xi}(G^{i})-\operatorname{deg}_{\xi}(\chi^{\vec{m}}))}$
@@
Now we discuss finiteness of bubbles, for fixed $\pi$, as it naturally arises in the previous theorem 2.4. A differential geometric counterpart is proven for the case when $\mathcal{X}_{t}(t\neq 0)$ are all smooth, compactifiable to metrized Kähler manifolds, in [Sun23, Corollary 1.12].
Theorem 2.12.
We fix a positive integer $n$. In the setup 1, 2, 3, we modify $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$. Then this modification process does not continue for infinitely many times for $\operatorname{dim}(X)=n$. Hence, we algebraically obtain a sequence of (weighted) algebraic bubbles $Z_{0},\cdots,Z_{2k}$ similarly to [Sun23, Cor 1.12].
(7) $=\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t)).$
for an arbitrary $0<|t|\ll 1$.
Proof.
The former half of Theorem 2.12 is a simple corollary of Theorem 2.4, that enhances its statements. The modification process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}_{\rm min}$ terminates by [XZ24]. We take $\mathcal{Y}$ as the end object of the repetition of the process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}$, towards deeper bubbles. Then the last equality (6) holds as otherwise we can repeat the process by Theorem 2.4, and the equality (7) is obvious from identification of $\mathcal{X}$ and $\mathcal{Y}$ outside fibers over $0\in\Delta$. ∎
Note that the last statement of the above Theorem 2.12 is a semistable reduction type theorem corresponding to virtual compactness of “moduli” of singularities with fixed normalized volumes. In particular, the following corollary to above is pointed out by M.Mauri. Our main reason of writing this is on our different method than usual resolution of singularities, rather than the result itself.
Corollary 2.13 (Semistable resolution).
Proof.
Theorem 2.14 (K-moduli of Fano cone [Od24a]).
Consider the set (resp., all $\mathbb{Q}$-Gorenstein families) of K-polystable (resp., K-semistable) $\mathbb{Q}$-Fano cone of $n$-dimension, with the fixed normalized volume $V$ of the vertices. Then, it forms a good proper (and separated) moduli algebraic space and its stacky enhancement, generalizing the K-moduli of $\mathbb{Q}$-Fano varieties.
2.2.3. Negative deformation

### v2

@@ 2.2.3. Details of the algebraic algorithmic construction
The K-semistability of $\mathcal{X}^{\prime}_{\rm min,0}$ in the sense of [Sun23, §5] follows from the construction as $W$ is a K-semistable log terminal cone ($\mathbb{Q}$-Fano cone) in the sense of [CS18, CS19]. Furthermore, from the positivity of the weights on the coordinates of $T^{\prime}$-action and the definition of elementary modification [AHLH23, §3], we conclude that the obtained test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ is a negative weight deformation in the sense of [CH22]111111just of rank $1$ i.e., with constant $\xi_{i}$ in their definition, [Od24b, §2]. Note that we do not know if any good choice of $T^{\prime}$ gives rise to a $\Theta$-strata in the original sense, and independence of $T^{\prime}$ is also a priori unclear (though it follows from [Od24b, 3.8 (2): canonicity claim]). These are why we need the higher rank setup and arguments in the generalized form [Od24b, Theorem 3.8] rather than the original in [AHLH23, §6].
@@
Now we are ready to discuss algebraic construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$, which follows a similar line of ideas as those for $\mathcal{X}^{\prime}_{\rm min}$, but also with some differences. We take a (non-canonical) algebraic subtorus $T^{\prime\prime}$ of $T$, with the rank $1$, which degenerates $W$ to the K-polystable degeneration as in [LWX21, Theorem 1.2]. Denote the corresponding base i.e., the partial compactification of $T^{\prime\prime}$ as $\overline{T}^{\prime\prime}$ which is isomorphic to $\mathbb{A}^{1}$. Recall that the construction in loc.cit is far from canonical. We still take one of them and denote the corresponding test configuration as $\mathcal{T}^{\prime\prime}\to\overline{T}^{\prime\prime}$. Henceforth we can and do assume $\mathcal{T}^{\prime\prime}$ is also achieved inside $M$, by enlarging $M$ inside the original base $\mathbb{A}_{k}^{p}$ if necessary, and $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) by retaking larger $l$, $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) and the embedding $\mathcal{X}\hookrightarrow\mathbb{C}^{l}\times\Delta$ (resp., $\mathcal{X}\hookrightarrow\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) is also realized in $M$ if necessary. Now we consider a couple of morphisms $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to M$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to M$) corresponding to $\mathcal{T}^{\prime}$ of degeneration of $X^{\prime}\rightsquigarrow W$, and corresponding to $\mathcal{T}^{\prime\prime}$ whose degeneration is of the form $W\rightsquigarrow C$.
@@
The above replacement $\mathcal{X}$ by $\mathcal{X}^{\prime}$ or $\mathcal{X}^{\prime\prime}$ or the existence of bubbles can first look in a sharp contrast to what happens in analogous global setup i.e., the separatedness of the moduli of log terminal projective Calabi-Yau varieties ([Od12, 4.2 (i), 4.3],  [^12] However, although 4.3 of op.cit is fine, there is an obvious typo in its 4.2(i), mixing up notation with canonically polarized case: in the left hand side $-K_{\mathcal{X}/\mathcal{C}}$ should be read $\mathcal{L}$ and we assume K-stability (log terminality) of the fibers of $\pi$. [Bou14, 2.1]).
@@
Our proof follows that of Theorem 2.4 closely, which essentially proves the case when $R=\mathbbm{k}[[t]]$ or $\mathbb{C}[[t]]^{\rm conv}$. For $\pi^{-1}(0)=\tilde{\mathcal{X}}_{0}$, we apply the same arguments as the beginning of the proof of Theorem 2.8. In particular, we take the same log terminal family $\mathcal{U}\subset\mathbb{A}_{\mathbbm{k}}^{l}\times\mathbb{A}_{\mathbbm{k}}^{m}$, which restricts to $\mathcal{V}\to M$ and consider the map $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to M\to\mathcal{M}$ which corresponds to $\tilde{\pi}$. As in the previous proof of Theorem 2.4, take the algebraic torus $T$ for $W$, the semistable tangent cone of $X$ and a rational polyhedral cone $\tau$ in the cocharacter lattice $N:={\rm Hom}(\mathbb{G}_{m},T)$ with $\xi\in\tau$ such that the base change of $\mathcal{V}$ over $U_{\tau}$ gives a flat isotrivial degeneration (over possibly higher dimensional base) with the central fiber $W$. If there is a uniform family of such isotrivial degenerations $X\rightsquigarrow W$ over a connected base, then from the same reason as in the proof of Theorem 2.4, i.e., by the use of index character ([MSY08, CS18, CS19]), the normalized volume are constant. Similarly as before, combined with the uniqueness of the $\widehat{\rm vol}$-minimizing valuation, this gives the canonical structure of higher $\Theta_{\tau}$-strata on $M_{c}$ for $c:=\widehat{\rm vol}(X,0)$. Then, we again apply the higher $\Theta$-stable reduction theorem [Od24b, §3.2, Theorem 3.8] to construct $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$. The rest of the constructions are almost the same, so we omit the essential repetition. ∎
@@ 2.2.4. Comparison with the differential geometric construction
Now, we discuss comparison with the differential geometric construction in [DS14, Sun23] and prove Theorem 2.11 and Theorem 2.13.
@@
We prove the item ($\mathrm{i}$) of Theorem 2.11. In this proof, we closely follow the arguments and notations of [Sun23] (cf., also [DS17]) by fitting it to our notation in Theorem 2.11. In the notation of loc.cit, suppose that his $X_{j}(j=1,2,\cdots)$ are taken as fibers of our $\pi$ i.e., $X_{j}=\overline{X}_{s_{j}}$ for some sequence $s_{j}\in\Delta$ with $s_{j}\to 0$ for $j=1,2,\cdots$ and $X_{\infty}=\pi^{-1}(0)=X$. Our $\xi$ corresponds to the real $N\times N$ diagonal matrix $\Lambda={\rm diag}(2^{\zeta_{1}},\cdots,2^{\zeta_{N}})$ with $\zeta_{i}\in\mathbb{R}_{>0}$s acting on the ambient space $\mathbb{C}^{N}$ in his notation. We crucially use the local holomorphic functions $G^{i},S^{i}_{j},L^{i}_{j}$ (and $H^{i}$, $H^{i}_{j}$ with higher degrees with respect to $\xi$) on the common ambient space $\mathbb{C}^{N}$ of the germs of $p_{j}\in X_{j}$s ($i$ was denoted as $b$ in loc.cit), $p_{\infty}\in X_{\infty}$ and the possible bubbles $Z$, the bounds of rescaling orders $\underline{k}$ and $\overline{k_{j}}$ as prepared in loc.cit. Then, we replace the sequence $\{j\}$ by its subsequence so that there is a sequence of real numbers $b_{j}(j=1,2,\cdots)$ with $b_{j}\to\infty$ and obtain
(8) $W_{1}:$ $=\lim_{j\to\infty}\Lambda^{b_{j}}\cdot X_{j}$
as a closed subscheme of $\mathbb{C}^{N}$. Let us write $\operatorname{deg}_{\xi}$ for the degree with respect to $\xi$ as in [Sun23, §4]. Here, each $L^{i}_{\infty}$ is obtained as a rescaled limit of $L^{i}_{j}$ for $j\to\infty$. Indeed, it can be achieved as follows. For each $i$, consider the set $L(i)$ of monomials $\chi^{\vec{m}}$ on the ambient space $\mathbb{C}^{N}$ whose $\xi$-degree is less than that of $G^{i}$. Its order is finite as $\zeta_{1},\cdots,\zeta_{N}$ are all positive. We denote the coefficient of $\chi^{\vec{m}}$ in $L^{i}_{j}$ as $c^{i}_{j}(\vec{m})$, then of course we have $c^{i}_{j}(\vec{m})\to 0\hskip 11.38109pt(j\to\infty)$ as $L^{i}_{j}\to 0\hskip 11.38109pt(j\to\infty)$. Now, for each fixed $j$, we consider
(9) $b_{j}:=\max_{i,\vec{m}\in L(i)}\dfrac{\log(|c^{i}_{j}(\vec{m})|)}{\log(2)\cdot(\operatorname{deg}_{\xi}(G^{i})-\operatorname{deg}_{\xi}(\chi^{\vec{m}}))}$
@@
Now we prove Theorem 2.13. The main point is confirmation of the compatibility of our construction in the proof of Theorem 2.4 with the case by case analysis in [dBS24], in each case.
proof of Theorem 2.13.
In all the three cases of Theorem 2.13, $X\ni\sigma(0)$ is locally an isolated singularity with local $\mathbb{C}^{*}$-action (with conical singularities) and hence in particular we can consider its finite dimensional smooth (unobstructed) deformation space ${\rm Def}(X\ni\sigma(0))$ of the singularity.
First, we consider Case ($\mathrm{i}$). In this case, there is a finite covering (the Weyl type covering) $\mathcal{D}:=\mathbb{C}^{k+1}\to{\rm Def}(X\ni\sigma(0))=\mathbb{C}^{k+1}/W\simeq\mathbb{C}^{k+1}$, where $W\subset S_{k+1}$ is the subgroup of the symmetric group of degree $k+1$, as the symmetry of the tuple $(\beta_{0},\cdots,\beta_{k})$, which acts by the permutation. The quotient is a Galois covering with the Galois group $G$. We consider the morphism $\Delta_{t}\to{\rm Def}(X\ni\sigma(0))$ which is associated to $\mathcal{X}(\supset\Delta_{t})\twoheadrightarrow\Delta_{t}$ and its lift $\Delta_{t}\to\mathcal{D}$ possibly after replacement of $\Delta_{t}$ by its finite covering (a component of the base change $\mathcal{D}\times_{{\rm Def}(X\ni\sigma(0))}\Delta_{t}$). We still denote the finite covering as $\Delta_{t}$ for simplicity, and denote as $\varphi\colon\Delta_{t}\to\mathcal{D}=\mathbb{C}^{k}$, $t\mapsto(s_{0}(t),\cdots,s_{k}(t))$ where each $s_{i}(t)$ is holomorphic with $s_{i}(0)=0$. In particular, the vanishing order of $s_{i}(t)$ at $t=0$ is positive which we denote as $d_{i}$. We set $d:=\min_{i}\{d_{i}\}$.
Note that the induced action of $\mathbb{C}^{*}$ as a subgroup of the automorphism group of $X\ni\sigma(0)$ on $\mathcal{D}$ is with the weights $(-1,\cdots,-1)$. Thus, from the proof of Theorem (2.4), $\mathcal{X}^{\prime}_{\rm min}$ naturally corresponds to
(10) $t\mapsto\varphi^{\prime}_{\rm min}(t):=(t^{-d}s_{0}(t),\cdots,t^{-d}s_{k}(t))\in\mathbb{C}^{k+1}=\mathcal{D}.$
Indeed, for $\epsilon>0$, we have that $\lim_{t\to 0}=[X\ni\sigma(0)=0]$, while $\varphi^{\prime}_{\rm min}(0)\neq$. On the other hand, it is easy to see that the flat (conical singular) Kähler metric on $\mathbb{C}$ is uniquely described as
$\prod_{0\leq i\leq k}|z-s_{i}(t)|^{\beta_{i}-1}|dz|$
(compare [dBS24, Lemma 1]), it follows that the differential geometric bubbling is obtained as the flat singular $\mathbb{C}$ given by $\varphi^{\prime}_{\rm min}(0)$. By replacing $\varphi$ by $\varphi^{\prime}_{\rm min}$ and iterative the procedure, we complete the proof of ($\mathrm{i}$) of Theorem 2.13.
The proof for the case ($\mathrm{ii}$) is more or less parallel to the case ($\mathrm{i}$) as above. We consider the well-known miniversal deformation space of $A_{k}$-singularity by Brieskorn-Grothendieck-Kas-Schlessinger-Slodowy. Recall it is also reconstructed as Kronheimer’s construction of a family of hyperKähler quotients [Kro89a], which we use. The deformation space can be taken as $\mathbb{C}^{k}$ and there is a natural Weyl cover $\mathcal{D}$ which kills the monodromy.
In our setup, due to the existence of ALE hyperKähler metrics $g_{t}$ with the $\partial\overline{\partial}$-exact Kähler form $\omega_{I,t}$, $\mathcal{X}\to\Delta_{t}$ can be written as (after base change to the Weyl cover i.e., take a component of $\Delta_{t}\times_{{\rm Def}(X\ni\sigma(0))}\mathcal{D}$)
$X_{t}=V(xy-\prod_{i=0}^{k}(z-s_{i}(t))\subset\mathbb{A}_{\mathbb{C}}^{3},$
where $s_{i}(t)$ is holomorphic. Indeed, each $X_{t}$ can be compared with the Gibbons-Hawking ansatz for appropriate $(0,s_{i}(t))\in\mathbb{R}\times\mathbb{C}\simeq\mathbb{R}^{3}(i=0,\cdots,k)$ with the harmonic potential
$\sum_{0\leq i\leq k}\frac{1}{|z-s_{i}(t)|}$
and the corresponding connection form as an essentially unique solution to the Bogomolony-type monopole equation, as explained in [LB91, p.302-303], [dBS24, §3.2, §3.3] and it follows from Kronheimer’s Torelli-type theorem [Kro89b] that they are isomorphic as hyperKähler varieties. It is obvious that $s_{i}(t)$ should be continuous but is even holomorphic by the second bulleted assumption on the holomorphicity of $\omega_{J,t}+\sqrt{-1}\omega_{K,t}$ on $\mathcal{X}\setminus X$. Now we can apply the discussion very similar to the proof of case ($\mathrm{i}$) i.e., consider the vanishing order of $s_{i}(t)$ at $t=0$ which is positive and denote by $d_{i}$ and set $d:=\min_{i}\{d_{i}\}$. Then we consider the map to the Weyl cover $\varphi^{\prime}_{\rm min}\colon\mathcal{D}(=\mathbb{C}^{k+1})\to{\rm Def}(A_{k})$ exactly as (10). Since the induced action of $\mathbb{C}^{*}$ as a subgroup of the automorphism group of $A_{k}$-singularity $X\ni\sigma(0)$ on $\mathcal{D}$ is again of the weights $(-1,\cdots,-1)$, we see that our algorithm in Theorem (2.4) gives $\mathcal{X}^{\prime}_{\rm min}$ which corresponds to $\varphi^{\prime}_{\rm min}$. Therefore, by comparing with [dBS24, §3.3, Theorem 3] and iterate the construction, we complete the proof.
The case ($\mathrm{iii}$) is easy as the deformation space ${\rm Def}(X=A_{1}\ni 0)$ is $1$-dimensional and smooth, which is realized as the $\mathbb{C}^{*}$-equivariat degeneration to the cone over quadric, which also serves as an example of (universal) negative weight deformation of Corollary 2.17 (a) in the next subsection. Each general fiber comes with the Stenzel metric [Ste93] for the complex quadric and it is easy to see that the construction in Theorem 2.4 gives the each general fiber. Thus, by [dBS24, §4.2 Proposition 1], we complete the proof. ∎
2.2.5. Negative weight deformation as scaling down